Three resistor of 6ohm , 10ohm and 15ohm are connected in parallel with a 6 volt battery find total resistance of the circuit ;total current ;current through the 10ohm resistor;potential difference across 15ohm resistor

Grade:10

1 Answers

Sushmita
6 years ago
1/R=1/6 +1/10 +1/15
1/R=5+3+3/30
1/R =10/30
R=3(total resistance)
 
V=IR
V=(6 given)
6=I×3
l=2 amp
Potential difference is same through all three resistors since connected in parallel